Calculate the force of attraction between a star of mass 20.00*10^30kg and the earth assuming the star is located 1.50*10^8km from the earth
1
Expert's answer
2020-04-20T10:15:40-0400

The Newton's law of gravitation says

"F=G\frac{m_1m_2}{r^2}\\
=6.67\times 10^{-11}\frac{20.00\times 10^{30}\times 5.97\times 10^{24}}{(1.50\times 10^{11})^2}\\
=3.54\times 10^{23}\:\rm N"